# Provably Unlearnable Examples

### v1

@@ 6. Provably Unlearnable Examples
6.1. Desiderata for Surrogates and PUEs
@@
As the surrogate for certification, $f_{\hat{\theta}}$ should exhibit a low classification error on ${\mathbb{D}}_{s}\oplus\delta$, signifying its successful capture of the UE distribution. Second, the recovery attack reveals that the manifold of UEs can be close to that of clean samples. Similar to the proof in Appendix F of the paper by Cohen et al. (Cohen et al., 2019), augmenting $f_{\hat{\theta}}$ with parametric noise maximizes the log-likelihood of correct predictions. This augmentation can tighten the certified $(q,\eta)$-Learnability.
@@
6.2. Random Weight Perturbation
@@
(13) $\min_{\hat{\theta}}$ $\ \mathbb{E}_{(x,y\sim{\mathbb{D}}_{s})}\ \min_{\delta,\ \delta^{\prime}\in\delta}\ \mathbb{E}_{\epsilon\sim\pi(0)}\ {\mathcal{L}}({f}_{\hat{\theta}+\epsilon}(x+\delta^{\prime}),y),$
$\ s.t.\ \forall\ \delta^{\prime}\in\delta,\ \|\delta^{\prime}\|\leq\xi,$
where $\delta^{\prime}$ is the corresponding noise added to $x$ and $\pi(0)$ is a train-time noise distribution with a mean of 00. In practice, we sample a set of $U$ random perturbations with a zero-mean Gaussian distribution, to train $\delta$ and $\hat{\theta}$. In the offline setting, we omit the inner minimization objective and only optimize $\hat{\theta}$. The training on the dataset ${\mathbb{D}}_{s}$ is thus to minimize the following empirical loss:
(14) $\min_{{\hat{\theta}}}\frac{1}{N}\sum_{j=1}^{N}\min_{\delta_{j},\|\delta_{j}\|\leq\xi}\frac{1}{U}\sum_{i=1}^{U}{\mathcal{L}}[f_{\hat{\theta}+\epsilon_{i}}(x_{j}+\delta_{j}),y_{j}],$
where $(x_{j},y_{j})$ is sampled from ${\mathbb{D}}_{s}$ and $\epsilon_{i}\sim\mathcal{N}(0,{\sigma^{\prime}}^{2}I)$ is the $i$-th noise added to $\hat{\theta}$. In each step of training, we gradually increase the value of $\sigma^{\prime}$ with a step size of $s$ to update $\hat{\theta}$ multiple times such that the classifier can better converge. The optimized perturbations $\delta$ are then superposed on ${\mathbb{D}}_{s}$ to make it unlearnable. The detailed algorithm for making the PUEs is depicted in Algorithm 3. More details of the optimization process are in Appendix A.
@@
Input: Surrogate classifier $\hat{f}_{\theta}$ with parameters $\theta$, training dataset ${\mathbb{D}}_{s}$, smoothing distribution $\mathcal{N}(0,\sigma^{2}I)$, $U_{train}$, $U_{perturb}$, noise cap $S$, step size $s$, training batch number $N$, train step number $M$, stop error rate $\tau$, validation set ${\mathbb{D}}^{\prime}_{s}$.
@@
1 Initialize:
2 $\delta\ \leftarrow\ [0]^{|{\mathbb{D}}_{s}|\times I}$
3 $error\leftarrow 1e^{9}$
4 while $error\geq\tau$ do
5       for $i\in 1,...,M$ do
6             Load a mini-batch $(x_{i},y_{i})$ from ${\mathbb{D}}_{s}$
@@
9             while $\sigma^{\prime}\leq S$ do
10                   $\sigma^{\prime}\ \leftarrow\ \sigma^{\prime}+s$
11                   for $j\in 1,...,U_{train}$ do
12                         $\epsilon_{j}\sim\ \mathcal{N}(0,\sigma^{\prime 2}I)$
13                         $l\ +=\ {\mathcal{L}}[f_{{\theta}+\epsilon_{j}}(x_{i}+\delta),y_{i}]$
14
15                  $\theta\ \leftarrow\ \textsc{Train\_step}(\theta,l/U_{train})$
16
17
18      for $i\in 1,...,N$ do
19             Load a mini-batch $(x_{i},y_{i})$ from ${\mathbb{D}}_{s}$
@@
22             while $\sigma^{\prime}\leq S$ do
23                   $\sigma^{\prime}\ \leftarrow\ \sigma^{\prime}+s$
24                   for $j\in 1,...,U_{perturb}$ do
25                         $\epsilon_{j}\sim\ \mathcal{N}(0,\sigma^{\prime 2}I)$
26                         $l\ +=\ {\mathcal{L}}[f_{{\theta}+\epsilon_{j}}(x_{i}+\delta),y_{i}]$
27
28                  $\delta\ \leftarrow\ \textsc{Opt\_step}(x_{i},\delta,l/U_{perturb})$
29
30
31      $error\ \leftarrow\ \textsc{Eval}(\theta,{\mathbb{D}}^{\prime}_{s}+\delta)$
32
[…]
@@ 7. Experiments
Baselines.  We consider both online surrogates and offline surrogates in the certification against CAs. We compare PUE with a tailored version of Error-Minimizing Noise (EMN) (Huang et al., 2021) in their certified $(q,\eta)$-Learnability to show the effectiveness of our online surrogate and the robustness of PUE. Moreover, we train offline surrogates on off-the-shelf PUE, EMN, and OPS (Wu et al., 2023) to compare their certified $(q,\eta)$-Learnability. When it comes to GAs, we conduct a series of empirical comparisons with EMN and OPS to show their robustness and clean test accuracy scores in cases where unauthorized classifiers escape from the certified parameter set $\hat{\Theta}$. Though our method does not restrict the type of noise applied to make data unlearnable (i.e., class-wise or sample-wise), we found that PUE surrogates trained with class-wise noises can converge rapidly while those trained by sample-wise noises do not converge very well. Therefore, in this section, we focus on class-wise noise in all experiments and comparisons.
@@
Training setups.  In each iteration of PUE optimization and online/offline surrogate training, the standard deviation of the noises added to the parameters is gradually increased from 00 with a step size of $0.05$ until the full noise level is reached. Furthermore, we start adding the weight perturbations after some warm-up steps to stabilize the training. Generally, we start perturbing weights after the classification error on UEs decreases below $50\%$. When generating UEs and fitting online surrogates, we set $M=10$ on CIFAR10, $M=20$ on CIFAR100, and $M=100$ on ImageNet. No random transformation is used in the surrogate training processes. In empirical experiments, we follow the previous training routines (Huang et al., 2021; Wu et al., 2023) and set the epoch to $60$ when training classifiers or launching recovery attacks on CIFAR10. For the training and recovery attacks on CIFAR100 and ImageNet, the epoch number is set to $100$. The weight decay rate is $5\times 10^{-4}$ on CIFAR10 and is $5\times 10^{-5}$ for CIFAR100/ImageNet. We apply an SGD with a momentum of $0.9$ and a learning rate of $0.1$ in all classifier/surrogate training runs.
Certification setups.  We select $0.25$ as the standard deviation of the train-time Gaussian noise to obtain the surrogates. During the certification process, we set $q=0.9$ and $\sigma\in\{0.25,0.8\}$ for all the surrogates. There are reasons behind this setting. First, $\sigma=0.25$ matches with the train-time noise, and can certify the highest $(q,\eta)$-Learnability at a fixed $\eta$ (see Section 7.3). Second, $\sigma=0.8$ is sufficient for certifying a parametric radius $\eta=1.0$ within which the recovery attack can restore most of the clean accuracy (refer to Figure 5 in Section 7.2). For each surrogate classifier, we sample $1000$ classifiers around it to compute the confidence interval of the QPS function under a confidence level of $0.99$. For each surrogate, we certify under various values of $\eta$ and record the certified learnability scores under each $\eta$. To compute the generalization $(q,\eta)$-Learnability, one can simply adding $\sqrt{\frac{1}{N}\log(\frac{2n}{\beta})}$ to the certified learnability scores given $n$, $N$ and $\beta$. For instance, for $n=1000$ and $\beta=0.01$, sampling a dataset with size $N=5000$ from the domain ${\mathcal{D}}$ as the test set to compute the generalization $(q,\eta)$-Learnability with attract a universal increase of $0.05$ to the certified learnability scores in the tables. Each run of certification on CIFAR10/CIFAR100 takes around one GPU hour on Nvidia Tesla P100. Certifying an ImageNet classifier requires 26.4 GPU hours. We apply offsets to the certified results to mitigate the impact of accuracy differences in the surrogates. More details of hyper-parameters and certification settings can be found in Appendix C.1.
7.1. Learnability Certification
@@
Figure 4. The clean test accuracy scores of surrogate classifiers gauged on clean CIFAR10 (left) and CIFAR100 (right) test set after recovery attacks. The best accuracy scores approximate the True Learnability of parameters inside the hypersphere centered at $\hat{\theta}$ with a radius of $\eta$.
@@
To make the comparisons, we introduce a baseline surrogate, PUE-B, which employs random weight perturbations in the training of the surrogate parameters but does not perturb the weights when optimizing the PAP noise. Therefore, PUE-B differs from EMN only in the training method of the surrogate and can be used as the baseline in the evaluation of PUEs. On the other hand, the surrogate trained without random weight perturbation is denoted as EMN. The comparison is made under two different levels of smoothing noise ($\sigma=0.25$ and $\sigma=0.8$). The results are presented in Table 1 and Table 2. According to Table 1, compared to EMN, PUE-B can be certified at a higher $(0.9,\eta)$-Learnability across all datasets. It can also be observed that the gain in the $(0.9,\eta)$-Learnability of PUE-B increases when the certified radius rises. On CIFAR100 and ImageNet, the gain becomes more obvious. The observation supports the effectiveness of the random weight perturbation strategy in reducing the gap between the $(q,\eta)$-Learnability and the True Learnability.
@@
Next, we compare the certified learnability of online surrogates trained by random weight perturbations on different UEs. We generated two versions of PUEs (i.e., PUE-1 and PUE-10) by setting the value of $U_{perturb}$ to $1$ and $10$, respectively. The certification results based on the corresponding online surrogates are recorded in Table 3 and Table 4. Generally, in all the tables, PUE-10 achieves the lowest $(0.9,\eta)$-Learnability. When the smoothing noise level matches the noise level of random weight perturbation, PUE-10 can diminish the $(0.9,\eta)$-Learnability on all datasets. However, when the parametric smoothing noise gets larger to $0.8$, PUE-10 has the best results at most of the certified radii across the three datasets but not for some cases at large $\eta$ values. This should be the consequence of the large smoothing noise used, which will be verified by the ablation study in Section 7.3. Saliently, all of the results on PUEs outperform that of PUE-B, meaning that optimizing the PAP noise over random weight perturbations can effectively suppress the $(0.9,\eta)$-Learnability.
@@
At last, we compare certification results based on offline surrogates trained on CIFAR10 and CIFAR100 with PAP noises. We find surrogates trained on CIFAR100 OPS cannot be augmented properly within the same training epochs used for PUE and EMN. We thus only compare with OPS on CIFAR10. The certified learnability scores are presented in Table 5 and Table 6. We have similar observations on the certification results based on online surrogates and offline surrogates. According to the results, PUEs certified based on offline surrogates outperform both EMN and OPS when $\sigma=0.25$. When $\sigma=0.8$, OPS performs better on CIFAR10 at small $\eta$. Similarly, EMN obtains slightly lower certified learnability scores at $\eta=0.7$. We suspect it is a consequence of the large smoothing noise, which leads to a loose certification. Additionally, the advantage of OPS can be a result of insufficient steps of training since classifiers usually require more epochs to converge on OPS. However, PUE still outperforms OPS when $\eta$ exceeds $0.4$.
@@
Compared to CIFAR10, CIFAR100 surrogates demonstrate lower best accuracy and recovered accuracy, which could be imputed to two possible reasons. First, CIFAR100 is by nature more difficult to learn than CIFAR10. Second, the CIFAR100 surrogate is trained by more iterations on the UEs and is harder to recover. However, it can be seen that the PUE surrogate still outperforms the EMN one on the tightness of certification. The EMN surrogate has the best accuracy ranging from $1.03\%$ to $13.69\%$ when $\eta$ increases, while the accuracy of the PUE surrogate can only be recovered to $1.40\%$ in the best case. The comparison shows the superiority of the PUE surrogate in certifying tighter $(q,\eta)$-Learnability. We also examine the impact of surrogate architecture towards the tightness of the certified $(q,\eta)$-Learnability. Please refer to Appendix C.2 for details.
Figure 5. The robustness of PUE and EMN against recovery attacks.
7.2. Robustness of PUE against GAs

### v2

@@ VI Provably Unlearnable Examples
VI-A Desiderata for Surrogates and PUEs
@@
As the surrogate for certification, $f_{\hat{\theta}}$ should exhibit a low classification error on ${\mathbb{D}}_{s}\oplus\delta$, signifying its successful capture of the UE distribution. Second, the recovery attack reveals that the manifold of UEs can be close to that of clean samples. Similar to the proof in Appendix F of the paper by Cohen et al. [21], augmenting $f_{\hat{\theta}}$ with parametric noise maximizes the log-likelihood of correct predictions. This augmentation can tighten the certified $(q,\eta)$-Learnability.
@@
VI-B Random Weight Perturbation
@@
$\min_{\hat{\theta}}$ $\ \mathbb{E}_{(x,y\sim{\mathbb{D}}_{s})}\ \min_{\delta^{\prime}\subseteq\delta}\ \mathbb{E}_{\epsilon\sim\pi(0)}\ {\mathcal{L}}({f}_{\hat{\theta}+\epsilon}(x+\delta^{\prime}),y),$ (13)
$\ s.t.\ \forall\ \delta^{\prime}\subseteq\delta,\ \|\delta^{\prime}\|\leq\xi,$
where $\delta^{\prime}$ is the corresponding noise added to $x$ and $\pi(0)$ is a train-time noise distribution with a mean of 00. In practice, we sample a set of $U$ random weight perturbations with a zero-mean Gaussian distribution, to train $\delta$ and $\hat{\theta}$. In the offline setting, we omit the inner minimization objective on $\delta$ and only optimize $\hat{\theta}$. The training on the dataset ${\mathbb{D}}_{s}$ is thus to minimize the following empirical loss:
$\min_{{\hat{\theta}}}\frac{1}{N}\sum_{j=1}^{N}\min_{\delta_{j},\|\delta_{j}\|\leq\xi}\frac{1}{U}\sum_{i=1}^{U}{\mathcal{L}}[f_{\hat{\theta}+\epsilon_{i}}(x_{j}+\delta_{j}),y_{j}],$ (14)
where $(x_{j},y_{j})$ is sampled from ${\mathbb{D}}_{s}$ and $\epsilon_{i}\sim\mathcal{N}(0,{\sigma^{\prime}}^{2}I)$ is the $i$-th noise added to $\hat{\theta}$. In each step of training, we gradually increase the value of $\sigma^{\prime}$ with a step size of $s$ to update $\hat{\theta}$ multiple times such that the classifier can better converge. The optimized perturbations $\delta$ are then superposed on ${\mathbb{D}}_{s}$ to make it unlearnable. The detailed algorithm for making the PUEs is depicted in Algorithm 3. More details of the optimization process are in Appendix -A.
@@
Input: Surrogate classifier $\hat{f}_{\theta}$ with parameters $\theta$, training dataset ${\mathbb{D}}_{s}$, smoothing distribution $\mathcal{N}(0,\sigma^{2}I)$, $U_{train}$, $U_{perturb}$, noise STD cap $S$, step size $s$, training batch number $N$, train step number $M$, stop error rate $\tau$, validation set ${\mathbb{D}}^{\prime}_{s}$.
@@
Initialize:
$\delta\ \leftarrow\ [0]^{|{\mathbb{D}}_{s}|\times d}$
$error\leftarrow 1e^{9}$
while $error\geq\tau$ do
       for $i\in 1,...,M$ do
             Load a mini-batch $(x_{i},y_{i})$ from ${\mathbb{D}}_{s}$
@@
             while $\sigma^{\prime}\leq S$ do
                   $\sigma^{\prime}\ \leftarrow\ \sigma^{\prime}+s$
                   for $j\in 1,...,U_{train}$ do
                         $\epsilon_{j}\sim\ \mathcal{N}(0,\sigma^{\prime 2}I)$
                         $l\ +=\ {\mathcal{L}}[f_{{\theta}+\epsilon_{j}}(x_{i}+\delta),y_{i}]$
                  $\theta\ \leftarrow\ \textsc{Train\_step}(\theta,l/U_{train})$
      for $i\in 1,...,N$ do
             Load a mini-batch $(x_{i},y_{i})$ from ${\mathbb{D}}_{s}$
@@
             while $\sigma^{\prime}\leq S$ do
                   $\sigma^{\prime}\ \leftarrow\ \sigma^{\prime}+s$
                   for $j\in 1,...,U_{perturb}$ do
                         $\epsilon_{j}\sim\ \mathcal{N}(0,\sigma^{\prime 2}I)$
                         $l\ +=\ {\mathcal{L}}[f_{{\theta}+\epsilon_{j}}(x_{i}+\delta),y_{i}]$
                  $\delta\ \leftarrow\ \textsc{Opt\_step}(x_{i},\delta,l/U_{perturb})$
      $error\ \leftarrow\ \textsc{Eval}(\theta,{\mathbb{D}}^{\prime}_{s}+\delta)$
$\hat{\theta}\ \leftarrow\ \theta$
@@ VII Experiments
Baselines.  We consider both online surrogates and offline surrogates in the certification against CAs. We compare PUE with a tailored version of Error-Minimizing Noise (EMN) [9] in their certified $(q,\eta)$-Learnability to show the effectiveness of our online surrogate and the robustness of PUE. Moreover, we train offline surrogates on off-the-shelf PUE, EMN, and OPS [35] to compare their certified $(q,\eta)$-Learnability. When it comes to GAs, we conduct a series of empirical comparisons with EMN and OPS to show their robustness and clean test accuracy scores in cases where unauthorized classifiers escape from the certified parameter set $\hat{\Theta}$. Though our method does not restrict the type of noise applied to make data unlearnable (i.e., class-wise or sample-wise), we found that PUE surrogates trained with class-wise noises can converge rapidly while those trained by sample-wise noises do not converge very well. Therefore, in this section, we focus on class-wise noise in all experiments and comparisons.
@@
Training setups.  In each iteration of PUE optimization and online/offline surrogate training, the standard deviation (STD) of the noises added to the parameters is gradually increased from 00 with a step size of $0.05$ until the full noise level is reached. Furthermore, we start adding the weight perturbations after some warm-up steps to stabilize the training. Generally, we start perturbing weights after the classification error on UEs decreases below $50\%$. When generating UEs and fitting online surrogates, we set $M=10$ on CIFAR10, $M=20$ on CIFAR100, and $M=100$ on ImageNet. No random data augmentation is used in the surrogate training processes. In empirical experiments, we follow the previous training routines [9, 35] and set the epoch to $60$ when training classifiers or launching recovery attacks on CIFAR10. For the training and recovery attacks on CIFAR100 and ImageNet, the epoch number is set to $100$. The weight decay rate is $5\times 10^{-4}$ on CIFAR10 and is $5\times 10^{-5}$ for CIFAR100/ImageNet. We apply an SGD with a momentum of $0.9$ and a learning rate of $0.1$ in all classifier/surrogate training runs.
Certification setups.  We select $0.25$ as the STD of the train-time Gaussian noise to obtain the surrogates. During the certification process, we set $q=0.9$ and $\sigma\in\{0.25,0.8\}$ for all the surrogates. There are reasons behind this setting. First, $\sigma=0.25$ matches with that of the train-time noise, and can certify the highest $(q,\eta)$-Learnability at a fixed $\eta$ (see Section VII-C). Second, $\sigma=0.8$ is sufficient for certifying a parametric radius $\eta=1.0$ within which the recovery attack can restore most of the clean accuracy (refer to Figure 5 in Section VII-B). For each surrogate classifier, we sample $1000$ classifiers around it to compute the confidence interval of the QPS function under a confidence level of $0.99$. For each surrogate, we certify under various values of $\eta$ and record the certified learnability scores under each $\eta$. To compute the generalization $(q,\eta)$-Learnability, one can simply adding $\sqrt{\frac{1}{N}\log(\frac{2n}{\beta})}$ to the certified learnability scores given $n$, $N$ and $\beta$. For instance, for $n=1000$ and $\beta=0.01$, sampling a dataset with size $N=5000$ from the domain ${\mathcal{D}}$ as the test set to compute the generalization $(q,\eta)$-Learnability will attract a universal increase of $0.05$ to the certified learnability scores in the tables. Each run of certification on CIFAR10/CIFAR100 takes around one GPU hour on Nvidia Tesla P100. Certifying an ImageNet classifier requires 26.4 GPU hours. We apply offsets to the certified results to mitigate the impact of accuracy differences in the surrogates. More details of hyper-parameters and certification settings can be found in Appendix -C.
VII-A Learnability Certification
@@
Figure 4: The clean test accuracy scores of surrogate classifiers gauged on clean CIFAR10 (left) and CIFAR100 (right) test set after recovery attacks. The best accuracy scores approximate the True Learnability of parameters inside the hypersphere centered at $\hat{\theta}$ with a radius of $\eta$.
@@
To make the comparisons, we introduce a baseline surrogate, PUE-B, which employs random weight perturbations in the training of the surrogate parameters but does not perturb the weights when optimizing the PAP noise. Therefore, PUE-B differs from EMN only in the training method of the surrogate and can be used as the baseline in the evaluation of PUEs. On the other hand, the surrogate trained without random weight perturbation is denoted as EMN. The comparison is made under two different levels of smoothing noise ($\sigma=0.25$ and $\sigma=0.8$). The results are presented in Table I and Table II. According to Table I, compared to EMN, PUE-B can be certified at a higher $(0.9,\eta)$-Learnability across all datasets. It can also be observed that the gain in the $(0.9,\eta)$-Learnability of PUE-B increases when the certified radius rises. On CIFAR100 and ImageNet, the gain becomes more obvious. The observation supports the effectiveness of the random weight perturbation strategy in reducing the gap between the $(q,\eta)$-Learnability and the True Learnability.
@@
Next, we compare the certified learnability of online surrogates trained by random weight perturbations on different UEs. We generated two versions of PUEs (i.e., PUE-1 and PUE-10) by setting the value of $U_{perturb}$ to $1$ and $10$, respectively. The certification results based on the corresponding online surrogates are recorded in Table III and Table IV. Generally, in all the tables, PUE-10 achieves the lowest $(0.9,\eta)$-Learnability. When the smoothing noise level matches the noise level of random weight perturbation, PUE-10 can diminish the $(0.9,\eta)$-Learnability on all datasets. However, when the parametric smoothing noise gets larger to $0.8$, PUE-10 has the best results at most of the certified radii across the three datasets but not for some cases at large $\eta$ values. This should be the consequence of the large smoothing noise used, which will be verified by the ablation study in Section VII-C. Saliently, all of the results on PUEs outperform that of PUE-B, meaning that optimizing the PAP noise over random weight perturbations can effectively suppress the $(0.9,\eta)$-Learnability.
@@
At last, we compare certification results based on offline surrogates trained on CIFAR10 and CIFAR100 with PAP noises. We find surrogates trained on CIFAR100 OPS cannot be augmented properly within the same training epochs used for PUE and EMN. We thus only compare with OPS on CIFAR10. The certified learnability scores are presented in Table V and Table VI. We have similar observations on the certification results based on online surrogates and offline surrogates. According to the results, PUEs certified based on offline surrogates outperform both EMN and OPS when $\sigma=0.25$. When $\sigma=0.8$, OPS performs better on CIFAR10 at small $\eta$. Similarly, EMN obtains slightly lower certified learnability scores at $\eta=0.7$. We suspect it is a consequence of the large smoothing noise, which leads to a loose certification. Additionally, the advantage of OPS can be a result of insufficient steps of training since classifiers usually require more epochs to converge on OPS. However, PUE still outperforms OPS when $\eta$ exceeds $0.4$.
@@
Compared to CIFAR10, CIFAR100 surrogates demonstrate lower best accuracy and recovered accuracy, which could be imputed to two possible reasons. First, CIFAR100 is by nature more difficult to learn than CIFAR10. Second, the CIFAR100 surrogate is trained by more iterations on the UEs and is harder to recover. However, it can be seen that the PUE surrogate still outperforms the EMN one on the tightness of certification. The EMN surrogate has the best accuracy ranging from $1.03\%$ to $13.69\%$ when $\eta$ increases, while the accuracy of the PUE surrogate can only be recovered to $1.40\%$ in the best case. The comparison shows the superiority of the PUE surrogate in certifying tighter $(q,\eta)$-Learnability. We also examine the impact of surrogate architecture towards the tightness of the certified $(q,\eta)$-Learnability. Please refer to Appendix -C for details. In addition, we check the validity of the certification under the stochasticity of different training runs to show that classifiers trained by adversaries can still fall inside the surrogate-certified $\hat{\Theta}$.
Figure 5: The robustness of PUE and EMN against recovery attacks.
Validity of the certification.  Recall that the validity of certified $(q,\eta)$-learnability is characterized by a certified parameter set. However, the training of classifiers faces stochasticity in weight initialization, mini-batch loading, stochastic gradient descent, etc. This implies that training on PUEs could lead to dissimilar model weights across different runs, so the PUEs that were certified by the surrogate classifier may not remain valid in subsequent training runs conducted by adversaries. To investigate the validity of PUEs for adversaries using mainstream training techniques, we train 10 ResNet-18 models on CIFAR10 PUEs with random weight initialization, batch loading orders, and data augmentations to check how diverse their trained weights can be. After training, we draw ten layers (nine convolution layers and one linear layer) and visualize the distribution of the converged weights. The visualization can be found along with our source code at https://github.com/Provably-Unlearnable-Examples/PUE. Furthermore, we calculate the pairwise difference for each parameter across the 10 classifiers and record the mean and STD of all the differences. According to the results, the weights from the 10 runs share similarities in their distributions, with a mean parameter difference of $-4.92\times 10^{-6}$ and an STD of $0.01$. Note that the certified parameter set $\hat{\Theta}$ is actually an infinite Gaussian mixture. For example, with $\sigma=0.25$ and $\eta=1.0$, $\hat{\Theta}$ has sufficiently large probability mass within $[\hat{\theta}-0.5,\hat{\theta}+0.5]$ for $\hat{\theta}$ with a dimensionality of $11$M (i.e., the rough parameter count of ResNet-18) and beyond, which suggests that the certified parameter set can cover classifier weights from different training runs with stochasticity. Therefore, $\hat{\Theta}$ certified by a surrogate $\hat{\theta}$ can effectively capture classifiers trained separately by adversaries following certain standard training procedures.
VII-B Robustness of PUE against GAs